3) Here’s a real-life example of how anatomy affects function. Domestic dogs have a diversity of skull

shapes ranging from those who look more like their wolf ancestors, such as huskies, to flat-faced

breeds, such as boxers. Because of these differences in morphology, two dogs of similar body mass

can have very different biomechanical forces in their jaws. The jaws below are representative of a

husky and a boxer (both can weigh approximately 25 Kg as adults). Points A and B are the canines and

molars, respectively. For both jaws, the muscle insertion point is on the bottom of the jaw and can

contract with a force of 3000N.

joint joint 3000N 3000N 15 cm 2 cm 1.5 cm 9.3 cm 0.7cm 1 cm

A) (+0.5 pt) What type of lever system best represents lower jaw of dogs? Type I, II or III?

B) Show your work for the calculations in parts i and ii and round answers to the nearest 10th.

i) (+0.5 pt) Using the information above, calculate the biting force for a husky at points A and B.

ii) (+0.5 pt) Now calculate the biting force for the boxer at points A and B.

iii) (+1 pt) Even though the jaw muscle contracts with similar force in both species (i.e. the same

Fi), each one has a different mechanical advantage and speed advantage. In your opinion, based

on jaw biomechanics, which would you want as a guard dog? Use either mechanical or speed

advantage (you should calculate it) to justify your opinion. (3 sentences max).

USE THE FOLLOWING FORMULAS:

F.L. = Fili Mechanical advantage: F/Fi = Li/L. Mechanical advantage: L/Li = Fi/F.
Answer #1

A) We can classify the dog jaw as a class 3 lever because the effort (by muscle near the lower jaw as shown in the figure and as described in the question) is applied between the load(teeth/food) and fulcrum(joint).

B)

1) For husky at point A

effort force Fi=3000N

effort arm length Li=1.5cm

load arm length Lo=(15+2+1.5)cm=18.5cm

load arm force Fo=?

FoLo=FiLi => Fo*18.5cm=3000N*1.5cm

Fo=243.243N=243.2N

For husky at point B

effort force Fi=3000N

effort arm length Li=1.5cm

load arm length Lo=(2+1.5)cm=3.5cm

load arm force Fo=?

FoLo=FiLi => Fo*3.5cm=3000N*1.5cm

Fo=1285.714N=1285.7N

2) For boxer at point A

effort force Fi=3000N

effort arm length Li=1cm

load arm length Lo=(9.3+0.7+1)cm=11cm

load arm force Fo=?

FoLo=FiLi => Fo*11cm=3000N*1cm

Fo=272.727N=272.7N

For boxer at point B

effort force Fi=3000N

effort arm length Li=1cm

load arm length Lo=(0.7+1)cm=1.7cm

load arm force Fo=?

FoLo=FiLi => Fo*1.7cm=3000N*1cm

Fo=1764.705N=1764.7N

3) Mechanical advantage =Fo/Fi

This measures the amplification of force that can be achieved by the lever. In an ideal case velocity ratio or speed advantage (Li/Lo) equals mechanical advantage. Note that the second formula for mechanical advantage is not correct in the question.

For husky at A = 243.2/3000=0.081

at B= 1285.7/3000=0.428

For boxer at A =272.7/3000=0.091

at B= 1746.7/3000=0.582

We see a boxer has a higher mechanical advantage at the canines and this is perfect for executing an attack of the intruder with greater force therefore the boxer has an edge over the husky as a guard dog. This means even though both jaw muscles produce the same force, the bite of a boxer has greater force because the jaw acts as a lever. The molars of the boxer also exert more force.